OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (anonymous):

v.w=3*(-4)+(-8)*(-2)+(-3)*(-6) =-12+16+18 =22